Solve the following equations:
(a/5) + 3 = 2...........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing 3 from LHS to RHS it becomes -3
Then,
= a/5 = 2–3
= a/5 = - 1
Now,
Multiply both side by 5,
= (a/5) × 5= -1 × 5
= a = -5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
a/5 + 3 = 2
a/5 = 2-3
a/5 = -1
a= -5
 
Hope it helps!
Thank you ✌️